# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v1

@@ 2.1 Relational Data Analysis
(1) End-to-End Reasoning: End-to-end reasoning requires the answer-generating LLM to directly generate the result with the provided prompt, which contains the time series data, instruction, and other context (e.g., news, events and logs). For example, SIGLLM alnegheimish2024can studies two paradigms for time series anomaly detection with LLMs. It mainly studies two paradigms: directly requesting LLMs for detection, or requesting them for prediction and then comparing the predicted results with actual data. TimeRAG yang2025timerag utilizes similar historical time series data to help the large model understand the current scene, achieving domain adaptation without a lot of training. It constructs a knowledge base with historical time series data, and retrieves reference sequence from it to strengthen the prediction. The reference sequences are selected by the distance measured by dynamic time warping (DTW) senin2008dynamic . Some work also studies the zero-shot or few-shot forecasting capability of LLMs in specific fields such as finance yu2023temporal and climate DBLP:journals/corr/abs-2411-13724 .
@@
(2) Agent-Based Methods: Agent-based methods utilize LLM agents to collaborate to complete analysis. Different agents often complete their own tasks under respective prompts. TimeCAP lee2025timecap is a framework with two agents and a multi-modal encoder. The summary agent generates a contextual summary of time series data, and the prediction agent uses the summary for prediction. The encoder generates predictions using both the data and summary. It also samples relevant text to reinforce the prediction agent. The final prediction is derived from the linear combination of the encoder prediction and the agent prediction. TimeXL jiang2025explainable proposes a more complex framework consisting of a multi-modal prototype-based encoder and three GPT-4o agents. The prototypes demonstrate the typical text patterns and temporal patterns. The encoder generates preliminary predictions and explanations based on time series and textual inputs. Three agents are respectively responsible for generating the prediction according to text input and explanations, comparing predicted and actual series to generate the feedback, improving text quality and triggering encoder retraining (e.g., updating the text input of the encoder based on feedback). The final prediction is derived from the weighted average of the encoder prediction and the agent prediction. The final prediction is also derived from the linear combination of the encoder prediction and the agent prediction. Considering the augmentation of prediction by relevant context, wang2024news selects relevant news to enhance forecasting by agents. A reasoning agent is instructed to select news as context of the prediction prompt, and a evaluation agent critiques the selected news and prediction result against the ground truth to update the logic of reasoning agent.
@@
$\bullet$ Time Series Alignment. Because of the modality differences between time series data and natural language, some methods consider aligning the modalities to better utilize LLMs to analyze time series data. These methods solves the challenge that LLMs may overlook temporal patterns, enabling them to better understand the time series format and analyze the dependencies. According to the mainly optimized architecture of models, they can be classified into designed encoder/decoder and fine-tuned LLM body. Their techniques are compared in the table I.
(1) Designed Encoder/Decoder: By designing and training special encoder or decoder modules, time series data is adapted to the input or output formats of LLMs with the backbone of LLMs unchanged. TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 proposes an effective embedding method to integrate time series data and textual context. It trains a patch reprogram module to encode time series data with text prototypes obtained by linearly probing the pre-trained word embeddings, and uses a frozen LLM embedder of Llama-7B to encode prompts as additional prefix. After the frozen LLM generates output embeddings, a trainable projection module is used to get the final forecasts.
To bridge the gap between structure, numerical features and language semantics, SEED li2025seed develops an encoding method with a token-aware structural encoder, a patch projection and alignment module and a semantic reprogramming module. It finally concatenates the tokens of multivariate time series data and task prompt as the input of LLM. Differently, TimeCMA liu2025timecma doesn’t put all the embeddings into LLM for the shortcoming of LLM-based solutions in learning disentangled embeddings. Instead, it encodes time series data into two branches, time series branch and LLM-empowered branch, combining them for multivariate time series forecasting via cross-modality alignment.
@@
(2) Fine-Tuned LLM Body: Based on designed encoder or decoder modules, some methods fine-tune LLM body on time-series tasks to adapt to the modalities of time series data, which further improves the LLMs’ ability to understand and analyze complex time series. CALF liu2025calf focuses on three aspects: embedding tokens, hidden features and prediction results. Accordingly, it applies the cross-modal match module that derives aligned text tokens queried by time tokens, the feature regularization loss between the textual and temporal hidden feature, and the output consistency loss between textual and temporal prediction. They reduce the distribution discrepancy in the corresponding aspect during the training process. S2IP-LLM pan2024sipllm , on the other hand, decomposes time series by seasonal-trend to explicitly encode each component. Tokens of components and relevant prompts are concatenated as input, while the component predictions are combined as the overall prediction. LLM4TS chang2025llm4ts proposes a two-stage fine-tuning strategy: The time series alignment stage aligns LLMs with time series data, while the forecasting fine-tuning stage fine-tunes LLMs on downstream forecasting tasks. In addition, a two-layer aggregation method integrates data of different time scales by pooling the sum of their embeddings, enhancing LLMs ’ ability to understand temporal information with specific time scale. Apart from prediction tasks, LLMFew chen2025large studies LLMs for multivariate time series classification. It introduces a patch-wise temporal convolution encoder (PTCEnc) with stacked causal convolution blocks for the alignment of time series representations and LLM’s input. And LoRA is used to fine-tune LLMs to improve their feature representation learning.
@@ 2.2 Graph Data Analysis
Different from relational data, graph data captures entities (vertices) and their interdependencies (edges) to explicitly model complex network semantics (e.g., social networks and knowledge graphs) beyond rigid tabular schemas. This structure is expressive, however, poses unique challenges, notably the vast search space and complex path reasoning involved in multi-hop queries defgraphdata . Compared with relational data analysis, graph data analysis involves more complex jobs like summarization based on the multi-hop relations across the graph vertices and reasoning over text-attributed graphs whose nodes and edges are associated with text liang2024natnl2gql ; zhou2024r3 . Graph data can not only be stored in relational databases, but also be stored and queried in knowledge graphs and accessed through SPARQL in RDF databases (e.g., Blazegraph blazegraph and GraphDB graphdb ) or Cypher in Neo4j neo4j .
@@
In contrast, the advent of LLMs offers transformative potential by leveraging their advanced reasoning capacities and cross-domain generalization abilities, which can (1) simplify the query writing costs (e.g., NL interfaces) and (2) achieve semantic-aware analysis unsupported in traditional ones.
@@
To align LLMs with graph databases in specific fields such as finance and medicine,  liang2024aligning proposes a pipleline combining the fine-tuning and the prompt method. NL-GQL pairs are generated by LLMs for fine-tuning, and the consistency is ensured by the background graph databases and two mutual verification self-instruct methods. In addition, due to the importance of relevant schema in the queries generating of LLMs, it will be extracted and integrated into prompts.
@@
2.3 Structured Data for LLM
Datasets for structured data, consisting of structured inputs, task definitions, and labels, play a crucial role in enhancing LLMs for structured data analysis. On one hand, training or fine-tuning with the training set is essential to adapt LLMs to domain-specific knowledge (e.g., finance, medical, climate) and structured data patterns (e.g., relational, time series, and graph). On the other hand, the test sets are widely used as benchmarks to evaluate the performance of LLMs.
The traditional method of dataset construction is usually the combination of real data collection and manual annotation. For example, the widely used table QA dataset WikiTQ pasupat2015compositional uses tables from Wikipedia and question-label pairs annotated by humans. Similarly, the NL2SQL dataset Spider yu2018spider uses databases collected from existing resources and question-SQL pairs annotated by college students. The correctness, diversity and readability of annotation results are ensured through multiple rounds of manual inspection. In addition, the correctness of SQL is further verified by script.
Relational Data Generation. REaLTabFormer solatorio2023realtabformer can generate relational datasets with the transformer. It first generates the parent table with a GPT-2 model and then generates the relational dataset conditioned on that with a Seq2Seq model. The GPT-2 and Seq2Seq model are trained on their corresponding task respectively. Another method hudovernik2024relational utilizes the structural representation capability of GNNs and the generative capability of diffusion models. It represents a relational database by a heterogeneous graph with the method in  xu2022synthetic , where a parent-child table pair connected by a foreign key is modeled as an attributed bipartite graph. Based on the GNN embeddings of the graph, the diffusion-based TabSyn  zhang2023mixed is used for conditional tabular data generation.
Graph Data Generation. To cope with the lack of large-scale graph structured datasets, a framework 10943225 to generate Large-Scale Graph Dataset is proposed. It generates structures and features separately by fitting the original graph, and aligns them to obtain the generated graph.  TKGGen presents the pipeline that generates temporal knowledge graph (TKG) datasets based on documents. It utilizes LLMs to generate timestamps for existing triplets, thereby obtaining the quadruples needed for constructing TKG.
@@ 3 LLM for Semi-Structured Data Analysis
Semi-structured data refers to data that are neither with strictly predefined schema like relational models nor raw data (e.g., plain text or images) abiteboul1997querying . Instead, they preserve partial organizational properties (e.g., tags, headers) and typically exhibit hierarchical or nested representation (e.g., County - Province - City in a nested JSON). This flexibility allows them to be represented in diverse formats, including web tables, spreadsheets, HTML, JSON, and XML. Broadly, such data can be categorized into markup languages and semi-structured tables, both of which pose challenges to downstream tasks due to their inherent structural complexity.
Takeaway In summary, semi-structured data remains difficult to handle due to irregular and nested structures.However, the strong semantic reasoning ability of LLMs shows great potential to overcome these challenges.Future research will focus on combining structure-aware methods with LLMs to achieve more robust and general solutions.
@@ 3.1 Markup Language
Markup languages (e.g., XML, JSON, HTML) are widely adopted for structuring and exchanging data. Nevertheless, research on integrating LLMs with markup languages remains relatively limited. Traditional methods for markup language understanding can be broadly classified into two categories. The first, Token Linearization, involves converting markup documents (e.g., HTML, JSON, Markdown) into plain text sequences and feeding them into PLMs herzig2020tapas . While straightforward, such methods are often too simplistic but hard to effectively capture the complexity of structured documents. To overcome this limitation, a second line of research, Tree/Graph-Based PLMs, seeks to explicitly encode structural information (e.g., DOM trees). However, owing to the limited generalization capacity and short context windows of PLMs, these models still struggle with issues such as the need for extensive pre-training, difficulties in handling large-scale data, and poor scalability.
Recently, LLM-based approaches have emerged as a promising alternative, offering stronger semantic reasoning and adaptability for handling the complexity of markup data. We conducted some research on this, and the work is shown in Figure 4.
@@
Markup Extraction. Markup Extraction refers to the task of extracting information from semi-structured data such as HTML and XML into the form of structured ones such as relational tables. Traditional approaches apply heuristic rules (e.g., find the content within the <h1><h1> or <title><title> tags wang2002detecting ) or data templates (e.g., extract the product name from <<h1 class=“product-name”>> crescenzi2001automatic ) to extract key information from markup information. However, these methods always face challenges such as template drift or high annotation costs. By leveraging the reasoning capabilities of LLMs, it becomes possible to directly extract and interpret hierarchical relationships, attributes, and nested structures from data, thus eliminating the reliance of the extraction task on index information such as tags or templates.
Evaporate arora2023language is a prototype system that leverages LLMs to derive structured views from semi-structured data. Given user-provided documents, it automatically generates structured representations of their contents. To support relational data extraction, Evaporate introduces the Evaporate-Code+ approach, which first employs LLMs to synthesize multiple candidate extraction functions. Specifically, the system generates functions through prompt engineering in two modes: one provides few-shot examples to the LLM and encourages the use of general-purpose Python functions, while the other specifies only the task requirements and directs the LLM to synthesize regular-expression–based solutions. Subsequently, a weak supervision–inspired aggregation mechanism is applied to assess and filter the candidate functions according to the quality of their outputs. Similarly, WebFormerwang2022webformer offers related functionality, but is tailored specifically to HTML by effectively incorporating its layout structure. Unlike Evaporate, WebFormer integrates website layout into transformer-based extraction by encoding DOM nodes as graph-attended HTML tokens and modeling text with efficient relative-position self-attention. Field tokens then guide heterogeneous cross-attention with HTML and text tokens, enabling accurate and scalable structured field extraction.
@@
Markup Information Querying. Markup information querying aims to extract user-intended information from data encoded in markup languages. While conceptually analogous to SQL querying over relational databases, it differs in that markup files are typically irregular and loosely structured, thereby introducing distinct processing challenges. To address these challenges, existing research has primarily explored tree-based approaches and NoSQL-based methods.
@@
$\bullet$ Tree-Based Approaches. XPath Agent li2024xpath adopts a two-stage pipeline for XPath query. In the Information Extraction (IE) stage, a lightweight LLM analyzes a simplified version of the webpage, where an HTML cleaning algorithm filters out irrelevant nodes and attributes, which finally generates a cleaned and compressed HTML. In the subsequent XPath Programming stage, a stronger LLM leverages the cleaned HTML and extracted information to construct XPath queries in a bottom-up manner. By progressively traversing from target nodes to parent elements and flexibly incorporating attributes such as class names and IDs.
@@
$\bullet$ NoSQL-Based Approaches. SMART lu2025bridging is a Text-to-NoSQL agent that adopts a multi-step framework integrating small language models (SLMs), retrieval-augmented generation (RAG), and LLMs. Specifically, it first employs an SLM to generate predicted schemas and an initial NoSQL query solely based on the natural language query. Next, it performs RAG over the dataset by leveraging the generated schema, the preliminary NoSQL query, and the original query. The retrieved results are then incorporated as few-shot exemplars and combined with the intermediate outputs to guide the LLM in producing a refined NoSQL query. The resulting query can subsequently be applied to large-scale semi-structured data.
@@
Markup Information Understanding. In contrast to querying markup languages, markup information understanding emphasizes leveraging LLMs to perform more complex operations on markup data (e.g., revision and error correction). To support such tasks, both end-to-end models and pipeline-based approaches have been developed for comprehensive understanding of markup files.
@@
DOM-LM deng2022dom proposes a structure-aware Transformer for modeling HTML documents by jointly encoding textual content and the DOM tree structure. To handle large DOM trees, it partitions them into manageable subtrees and incorporates structure-aware positional embeddings (e.g., node depth, parent indices, and sibling indices) to capture hierarchical context. The model is further trained with a self-supervised masking strategy applied across both tokens and DOM nodes. In contrast, MarkupLM li2021markuplm extends token representations with four complementary perspectives. In addition to text embeddings for encoding document words, 1D positional embeddings for capturing horizontal order, and segment embeddings for distinguishing paragraph information, it introduces XPath embeddings. The method first decomposes an XPath into hierarchical layers consisting of tag names and subscripts. For each layer, the embeddings of the tag and its corresponding subscript are combined to form a unit-level embedding. These embeddings across all layers are then concatenated and passed through a feed-forward network (FFN) to generate the final XPath representation. For pre-training, MarkupLM employs three markup-specific objectives: (1) Masked Markup Language Modeling (MMLM) to jointly capture textual and structural information, (2) Node Relationship Prediction (NRP) to learn hierarchical relations among DOM nodes, and (3) Title–Page Matching (TPM) to assess global document coherence.
@@
WebLM xu2024hierarchical introduces a hierarchical multimodal pre-training framework for webpage understanding. Specifically designed for HTML, it integrates structural information, textual content, and webpage screenshots within a unified Transformer encoder. The framework first aligns visual features with their corresponding HTML nodes and then employs two novel training objectives: (1) Tree Structure Prediction (TSP), which predicts parent–child and sibling relations among HTML nodes to strengthen structural understanding, and (2) Visual Misalignment Detection (VMD), which perturbs the size of node-associated visual regions to improve robustness against layout variations.
@@
Insights of Markup Language Analysis. Currently, most studies on markup language understanding continue to rely on pre-trained language models (PLMs) for two main reasons. (1) Controllability: PLMs offer a fully controllable environment, where model size, training data, and task-specific settings can be explicitly defined. In contrast, LLM-based APIs are typically closed-source and involve high costs for fine-tuning and inference. (2) Structural Adaptation: PLMs can be tailored with specialized tokenization and positional encodings to capture the hierarchical structure of HTML/XML deng2022dom ; xu2024hierarchical . However, general-purpose LLMs often depend on prompt engineering, resulting in redundant tokenization and excessively long prompts, while also struggling to jointly model semantic and structural information (e.g., converting an HTML table into plain text requires an auxiliary schema to record the layout of each cell).
@@
$\bullet$ Information Compression. Information compressiong aims to address the challenges of encoding markup languages by providing LLMs with more concise representations. Traditional approaches, such as structural prompting or schema-aware prompting, often result in redundant rules and excessively long prompts. In contrast, segment-wise prompting partitions long documents into manageable segments and selectively filters useful paragraph information. For renderable languages such as HTML or LaTeX, vision–language models may further assist in file partitioning and alignment with encoded content. For instance, using the VLM to help with visual anchor extraction, and generating a tree-structure to encode the HTML or LaTeX document. Then, we can flexibly select the anchor information according to the task and give it to LLM, which also reduces prompt length and retains the necessary structure information.
$\bullet$ Tool Integration. Tool Integration for LLM is an insight inspired by LLM Agent. This approach is to enable LLM with abilities to change the outer environment. There are already existing methods such as DSL-Based Soft Tools, Retrieval-based Code Generation, and API Calling frameworks. However, those tools are usually a little rigid, like API Calling frameworks is a fixed mapping between natural language and predefined functions.So we need to explore more dynamic and adaptive approaches, such as integrating planning modules, learning-based tool selection, and even self-reflective reasoning loops.

### v2

@@ 2.1 Relational Data Analysis
(1) End-to-End Reasoning: End-to-end reasoning requires the LLM to generate results directly from a prompt containing time series data, instructions, and contextual information (e.g., news, events, logs). For instance, SIGLLM alnegheimish2024can explores anomaly detection by either directly querying the LLM or using it for prediction and comparing results with actual data. TimeRAG yang2025timerag leverages historical time series for domain adaptation, retrieving reference sequences via dynamic time warping (DTW) senin2008dynamic to strengthen predictions. Other studies examine LLMs’ zero- or few-shot forecasting abilities in domains like finance yu2023temporal and climate DBLP:journals/corr/abs-2411-13724 .
@@
(2) Agent-Based Methods: Agent-based methods employ multiple LLM agents, each completing tasks under specific prompts, to collaboratively perform analysis. TimeCAP lee2025timecap uses two agents and a multi-modal encoder. The summary agent generates contextual summaries of time series data, while the prediction agent uses these summaries for forecasting. The encoder integrates both data and summaries, reinforced by sampled relevant text, and the final prediction derived from the linear combination of encoder and agent outputs. TimeXL jiang2025explainable extends this with a prototype-based encoder and three agents based on GPT-4o. The encoder produces preliminary predictions and explanations based on time series and textual inputs, while one agent generates predictions from text and explanations, another compares predictions with actual series to provide feedback, and the third refines text quality and triggers encoder updates. The final prediction is a weighted combination of encoder and agent outputs. For context-augmented forecasting wang2024news , a reasoning agent selects relevant news to provide context for predictions, while an evaluation agent critiques both the news selection and prediction against ground truth, updating the reasoning agent’s logic.
@@
$\bullet$ Time Series Alignment. To address modality differences between time series data and natural language, some methods align the modalities to enable LLMs to capture temporal patterns and dependencies more effectively. Based on the model architecture, these approaches are classified into designed encoder/decoder frameworks and fine-tuned LLM bodies, with a comparison summarized in Table I.
(1) Designed Encoder/Decoder: Specialized encoder or decoder modules adapt time series data to LLM input or output formats without modifying the LLM backbone. TIME-LLM DBLP:conf/iclr/0005WMCZSCLLPW24 integrates time series and textual context by encoding time series with a patch reprogram module using text prototypes from pre-trained embeddings, prepending prompts encoded by a frozen LLaMA-7B embedder, and generating forecasts through a trainable projection module. SEED li2025seed bridges structural, numerical, and semantic gaps using a token-aware structural encoder, patch projection and alignment, and semantic reprogramming, concatenating multivariate time series tokens with the task prompt as LLM input. In contrast, TimeCMA liu2025timecma avoids feeding all embeddings into the LLM due to its limitations in learning disentangled representations. It encodes time series through two branches, a time series branch and an LLM-empowered branch, and combines them via cross-modality alignment for multivariate forecasting.
@@
(2) Fine-Tuned LLM Body: Some methods fine-tune LLMs with encoder or decoder modules on time-series tasks, enhancing their understanding and analysis of complex time series. CALF liu2025calf targets embeddings, hidden features, and predictions, applying cross-modal token alignment to derive aligned text tokens queried by time tokens, feature regularization loss between the textual and temporal hidden feature and output consistency loss between textual and temporal prediction to reduce distribution discrepancies during training. S2IP-LLM pan2024sipllm , on the other hand, decomposes time series by seasonal-trend to explicitly encode each component. Tokens of components and relevant prompts are concatenated as input, while the component predictions are combined as the overall prediction. LLM4TS chang2025llm4ts uses a two-stage fine-tuning which aligns LLMs with time series data, then fine-tuning on forecasting tasks. A two-layer aggregation pools the sum of embeddings across time scales to improving temporal understanding ability. Apart from prediction tasks, LLMFew chen2025large addresses multivariate time series classification, using a patch-wise temporal convolution encoder (PTCEnc) to align time series with LLM input, and fine-tunes LLMs via LoRA to enhance feature representation.
@@
Relational Data Generation.
On one hand, training or fine-tuning with the training set is essential to adapt LLMs to domain-specific knowledge (e.g., finance, medical, climate) and structured data patterns (e.g., relational, time series, and graph). On the other hand, the test sets are widely used as benchmarks to evaluate the performance of LLMs.
The traditional method of dataset construction is usually the combination of real data collection and manual annotation. For example, the widely used table QA dataset WikiTQ pasupat2015compositional uses tables from Wikipedia and question-label pairs annotated by humans. Similarly, the NL2SQL dataset Spider yu2018spider uses databases collected from existing resources and question-SQL pairs annotated by college students. The correctness, diversity and readability of annotation results are ensured through multiple rounds of manual inspection. In addition, the correctness of SQL is further verified by script. MMTU xing2025mmtu provides a comprehensive collection of 52 such manually curated benchmarks on structured tables, organized across different tasks and benchmark datasets.
REaLTabFormer solatorio2023realtabformer can generate relational datasets with the transformer. It first generates the parent table with a GPT-2 model and then generates the relational dataset conditioned on that with a Seq2Seq model. The GPT-2 and Seq2Seq model are trained on their corresponding task respectively. Another method hudovernik2024relational utilizes the structural representation capability of GNNs and the generative capability of diffusion models. It represents a relational database by a heterogeneous graph with the method in  xu2022synthetic , where a parent-child table pair connected by a foreign key is modeled as an attributed bipartite graph. Based on the GNN embeddings of the graph, the diffusion-based TabSyn  zhang2023mixed is used for conditional tabular data generation.
@@ 2.2 Graph Data Analysis
Unlike relational data, graph data captures entities (vertices) and their relationships (edges), enabling explicit modeling of complex network semantics (e.g., social networks, knowledge graphs). While highly expressive, this structure introduces unique challenges, including vast search spaces and intricate multi-hop reasoning defgraphdata .
Graph data analysis, compared with relational data, entails more complex tasks, such as summarizing multi-hop relationships and reasoning over text-attributed graphs, where nodes and edges carry textual information liang2024natnl2gql ; zhou2024r3 . Graphs can be stored not only in relational databases but also in knowledge graphs, accessed via SPARQL in RDF databases (e.g., Blazegraph blazegraph , GraphDB graphdb ) or Cypher in Neo4j neo4j .
@@
In contrast, the advent of LLMs offers transformative potential by leveraging their advanced reasoning capacities and cross-domain generalization abilities, which can $(i)$ simplify the query writing costs (e.g., NL interfaces) and $(ii)$ achieve semantic-aware analysis unsupported in traditional ones.
@@
To align LLMs with graph databases in specific fields such as finance and medicine,  liang2024aligning proposes a pipeline combining the fine-tuning and the prompt method. NL-GQL pairs are generated by LLMs for fine-tuning, and the consistency is ensured by the background graph databases and two mutual verification self-instruct methods. In addition, due to the importance of relevant schema in the queries generating of LLMs, it will be extracted and integrated into prompts.
@@
Graph Data Generation. To cope with the lack of large-scale structured graph datasets, a framework 10943225 to generate Large-Scale Graph Dataset is proposed. It generates structures and features separately by fitting the original graphs, and aligns them to obtain the generated graph.  TKGGen presents the pipeline that generates temporal knowledge graph (TKG) datasets based on documents. It utilizes LLMs to generate timestamps for existing triplets, thereby obtaining the quadruples needed for constructing TKG.
Takeaway: For structured data analysis, LLM-based agents demonstrate significant abilities: Complex Data Understanding (O1): LLMs capture structural and semantic relations across tables, graphs, and time series data, enabling unified understanding. NL-Based Interface (O2): LLMs translate user intents into executable queries (e.g., SQL, graph queries), facilitating intuitive data interaction. Semantic Operators (O3): LLMs integrate symbolic and neural reasoning to execute compositional operations with semantic awareness. Autonomous Evolution (O4): Through continual learning and schema adaptation, LLM agents refine their analytical abilities and evolve with dynamic data and tasks.
@@ 3 LLM for Semi-Structured Data Analysis
Semi-structured data are neither strictly schema-defined like relational data nor completely raw like text or images abiteboul1997querying . They retain partial organizational properties (e.g., tags, headers) and often have hierarchical or nested structures (e.g., CountyProvinceCity in JSON). This allows representation in diverse formats such as web tables, spreadsheets, HTML, JSON, and XML. They can be broadly categorized into markup languages and semi-structured tables, both challenging downstream tasks due to their structural complexity.
@@ 3.1 Markup Language
Markup languages (e.g., XML, JSON, HTML) are widely used for structuring and exchanging data, yet research on integrating LLMs with them remains limited. Traditional approaches for markup understanding fall into two categories. Token Linearization converts markup documents into plain text sequences for PLMs herzig2020tapas , but this often oversimplifies and fails to capture structural complexity. Tree/Graph-Based PLMs explicitly encode structures (e.g., DOM trees), yet they suffer from limited generalization, short context windows, high pre-training costs, and poor scalability.
LLM-based methods have recently emerged as a promising alternative, offering stronger semantic reasoning and greater adaptability for complex markup data, which is summarized in Figure 4.
@@
Markup Extraction. Markup Extraction aims to transform semi-structured data (e.g., HTML, XML) into structured formats such as relational tables. Traditional methods rely on heuristic rules (e.g., extracting content within <h1><h1> or <title><title> tags wang2002detecting ) or data templates (e.g., extracting product names from <<h1 class=“product-name”>> crescenzi2001automatic ), but they often suffer from template drift and high annotation costs. Leveraging LLMs enables direct extraction and interpretation of hierarchical relationships, attributes, and nested structures, reducing dependence on tags or templates.
Evaporate arora2023language is a prototype system that uses LLMs to generate structured views from semi-structured documents. It employs Evaporate-Code+, which first synthesizes multiple candidate extraction functions via LLMs. Functions are generated in two modes, where one provides few-shot examples encouraging general-purpose Python solutions, and the other specifies only task requirements to produce regex-based solutions. A weak supervision–inspired aggregation mechanism then evaluates and filters candidates based on output quality.
Similarly, WebFormer wang2022webformer targets HTML specifically by leveraging its layout structure. It encodes DOM nodes as graph-attended tokens and models text with relative-position self-attention. Field tokens guide heterogeneous cross-attention with HTML and text tokens, enabling accurate and scalable structured field extraction.
@@
Markup Information Querying. Markup information querying extracts user-intended data from markup languages. While conceptually similar to SQL querying over relational databases, markup files are often irregular and loosely structured, posing unique challenges. Existing research mainly explores tree-based and NoSQL-based approaches.
@@
$\bullet$ Tree-Based Approaches. XPath Agent li2024xpath uses a two-stage pipeline for XPath queries. In the information extraction stage, a lightweight LLM analyzes a simplified webpage, with irrelevant nodes and attributes removed via HTML cleaning algorithm to produce a compressed HTML. In the XPath Programming stage, a stronger LLM constructs XPath queries bottom-up, traversing from target nodes to parent elements and flexibly using attributes like class names and IDs.
@@
$\bullet$ NoSQL-Based Approaches. SMART lu2025bridging is a Text-to-NoSQL agent that integrates small language models (SLMs), retrieval-augmented generation (RAG), and LLMs. An SLM first predicts schemas and an initial NoSQL query from the natural language input. RAG then retrieves relevant data using the preliminary query and schema, which, combined with few-shot exemplars, guides the LLM to generate a refined NoSQL query applicable to large-scale semi-structured data.
@@
Markup Information Understanding. Unlike markup querying, markup information understanding focuses on using LLMs for more complex operations on markup data (e.g., revision and error correction). To support this, both end-to-end and pipeline-based models have been developed for comprehensive markup file understanding.
@@
DOM-LM deng2022dom introduces a structure-aware Transformer that jointly encodes HTML text and DOM tree structure. To handle large DOM trees, it partitions them into subtrees and uses structure-aware positional embeddings (e.g., node depth, parent, and sibling indices) to capture hierarchy. The model is trained with a self-supervised masking strategy over both tokens and DOM nodes. MarkupLM li2021markuplm extends token representations from four perspectives, including text embeddings, 1D positional embeddings for horizontal order, segment embeddings for paragraph distinctions, and XPath embeddings. XPath embeddings are created by decomposing an XPath into hierarchical layers of tags and subscripts, combining each layer’s tag and subscript embeddings, concatenating across layers, and passing through a feed-forward network to produce the final representation. For pre-training, MarkupLM uses three markup-specific objectives. Masked Markup Language Modeling (MMLM) for joint textual-structural understanding, Node Relationship Prediction (NRP) to learn DOM hierarchies, and Title–Page Matching (TPM) for global document coherence.
@@
WebLM xu2024hierarchical presents a hierarchical multimodal pre-training framework for HTML webpage understanding, integrating structural, textual, and visual information within a unified Transformer. Visual features are aligned with HTML nodes, and two training objectives are used. Tree Structure Prediction (TSP) captures parent–child and sibling relations, and Visual Misalignment Detection (VMD) improves robustness to layout variations.
@@
Insights of Markup Language Analysis. Most studies on markup language understanding still rely on PLMs for two reasons. $(i)$ Controllability: PLMs allow explicit control over model size, training data, and task settings, whereas LLM APIs are often closed-source and costly. $(ii)$ Structural Adaptation: PLMs can use specialized tokenization and positional encodings to capture HTML/XML hierarchies deng2022dom ; xu2024hierarchical , while general-purpose LLMs depend on prompt engineering, leading to redundant tokenization, long prompts, and difficulty jointly modeling semantics and structure (e.g., converting an HTML table to text requires auxiliary schemas).
@@
$\bullet$ Information Compression. Information compression addresses the challenge of encoding markup languages by providing LLMs with more concise representations. Traditional methods, such as structural or schema-aware prompting, often produce redundant rules and long prompts. Segment-wise prompting partitions documents into manageable segments and selectively filters relevant content. For renderable languages like HTML or LaTeX, vision–language models can assist by extracting visual anchors and generating tree-structured encodings, allowing LLMs to receive task-relevant information while reducing prompt length and preserving structural context.
$\bullet$ Tool Integration. Tool Integration for LLMs aims to enable LLMs to interact with and modify external environments. Existing methods such as DSL-based soft tools, retrieval-based code generation, and API-calling frameworks, are often rigid, with APIs providing fixed mappings from natural language to predefined functions.More dynamic approaches are needed, including planning modules, learning-based tool selection, and self-reflective reasoning loops.